OpenStudy (anonymous):

Five of Ethan's family members went on a 4 and one-sixth hour road trip. If they all shared the drive time equally, how many hours did each family member drive? five-sixths 1 and seven-twenty-fourths twenty-three thirtieths 1 and one-fifth

OpenStudy (anonymous):

\[4\tfrac{1}{6}\div 5\]

OpenStudy (anonymous):

to divide turn \(4\tfrac{1}{6}\) in to an improper faction do you know how to do that?

OpenStudy (anonymous):

"no" is a fine answer, i am just asking

OpenStudy (anonymous):

\[4\tfrac{1}{6}=\frac{4\times 6+1}{6}=\frac{25}{6}\] and so \[\frac{25}{6}\div 5=\frac{25}{6}\times \frac{1}{5}=\frac{5}{6}\]